# Prioritization First, Principles Second: An Adaptive Interpretation of Helpful, Honest, and Harmless Principles

### v4

@@ 7 Open Challenges
While the proposed reference framework offers an adaptable approach for interpreting the HHH principle, it is subject to several limitations that merit consideration. First, the framework does not adequately address fundamental challenges inherent to the HHH principle themselves. For example, it provides limited support for mitigating issues like out-of-distribution (OOD) adversarial attacks (Zou et al., 2023; Huang et al., 2024c, 2025), where models encounter inputs significantly deviating from their training data. Such scenarios often require specialized robustness and generalization methods for enhancement, which fall outside the scope of this framework.
@@
Additionally, the framework faces implementation challenges–inherent difficulties associated with multi-objective optimization (MOO). Balancing helpfulness, honesty, and harmlessness requires trade-offs that are complicated by normative assumptions embedded in the optimization objectives. The lack of precise quantification methods for the interaction between these principles complicates achieving a theoretically optimal balance, limiting the framework’s applicability in certain high-stakes or ambiguous scenarios. Following up recent studies propose potential methods for MOO alignment (Bai et al., 2022a; Mukherjee et al., 2024), more efforts should be made in addressing this challenge.
@@
Moreover, a fundamental question remains: Can human values, ethics, and honesty be embedded into AI systems with sufficient certainty? This challenge is an requirement and pivotal for the proposed framework to function effectively, yet it remains an open problem (Sorensen et al., 2024). Human values are inherently dynamic, context-dependent, and often subjective, making it difficult to codify them in a way that guarantees reliable and universally accepted AI behavior (Scherrer et al., 2024). Furthermore, ethical principles may conflict in ambiguous scenarios, requiring nuanced decision-making that AI models currently struggle to replicate (e.g., (Moore et al., 2024; Shrivastava et al., 2024)). Without reliable mechanisms to encode these values, the framework risks being constrained by the same uncertainties that challenge broader AI alignment efforts.
@@
Notably, the criticism that HHH principles are merely high-level heuristics and lack operational substance neglects two important facts. First, the original intention of HHH was to anchor alignment efforts in human-centered values, offering interpretable and ethically grounded objectives for training and evaluation. While it is true that these principles are polysemantic and context-sensitive-as are any high-level ethical ideals such as autonomy or respect-this does not diminish their value. Rather, it reinforces the necessity of the adaptive framework we propose, which systematically resolves ambiguities through contextual modeling, priority ordering, and risk assessment. Second, asserting that operational guidance has moved beyond HHH (e.g., to model specs or rule-based reward shaping) overlooks the fact that these newer approaches often implicitly instantiate the HHH values. The 30-40 page model specs (OpenAI, 2024a) and deliberative alignment strategies do not replace HHH; they concretize it. Our contribution lies in formalizing this process: instead of discarding HHH for being abstract, we operationalize it adaptively-preserving its normative clarity while enabling contextual application.
@@ Appendix A Model, Framework and Benchmark Example
HHH (Askell et al., 2021) ˇ ˇ
TrustLLM (Huang et al., 2024a) ˇ ˚
NIST (NIST, 2024) ˇ ˚
AI Safety and Security (Qi et al., 2024) ˇ ˘
Risk Assessment (Kapoor et al., 2024) ˇ ˘
Transparency (Bommasani et al., 2024) ˇ ˘
Table 2: Typical principles or frameworks for guiding AI model alignment (excluding government laws or acts, e.g., EU AI Act (EU, 2024b) and AI Bill of Rights (House, 2022)). ˇ, ˚, and ˘ indicate different emphasis on trustworthiness or utility.
@@
Llama 2 (Touvron et al., 2023b) LLM 2023
LLaVA (Liu et al., 2024b) VLM 2023
Qwen (Bai et al., 2023) LLM 2023
InternLM (Team, 2023) LLM 2023
Llava-med (Li et al., 2024a) VLM 2024
Aya model (Üstün et al., 2024) LLM 2024
LLaVA-Plus (Liu et al., 2025) VLM 2024
MiniCPM (Hu et al., 2024) LLM 2024
@@
MTBench (Zheng et al., 2023)
MMLU (Hendrycks et al., 2020)
HumanEval (Chen et al., 2021)
Libra-Leaderboard (Li et al., 2024b)
DecodingTrust (Wang et al., 2023)
TrustLLM (Huang et al., 2024b)
HonestLLM (Gao et al., 2024)
BeHonest (Chern et al., 2024)
@@ Appendix B Example of Priority Order
  • Healthcare: Honesty is uttermost critical, as incorrect diagnoses or misleading AI outputs can directly endanger lives. Harmlessness safeguards sensitive medical records, ensuring AI maintains their privacy yet with accurate responses Liu et al. (2023b). Helpfulness is desirable but meaningless if accuracy is compromised.
@@ Appendix C Case Study of the 3H principles for developing Chemistry Foundation Models
Alignment Auditing Context-Specific Benchmarks: Design benchmarks that align with the prioritized dimensions. For instance: 1). Measured by the model’s predictive accuracy and relevance to user tasks (Guo et al., 2023), 2). Evaluated through safety checks on generated molecular outputs. 3). Assessed by comparing generated outputs with verified scientific data (Guo et al., 2024a; Bushuiev et al., 2024). Multi-Dimensional Integration: Use complementary benchmarks to capture performance holistically across different dimensions.
Governance Infrastructure Transparency: In the chemistry domain, models must provide interpretable outputs, highlighting key chemical features influencing predictions and annotating data sources such as PubChem (Kim et al., 2021) or ChEMBL (Mendez et al., 2019). Confidence scores and error margins should accompany predictions to ensure reliability. Governance: It requires strict validation of outputs against experimental data, safeguards against misuse (e.g., toxic compound generation), and compliance with safety standards like EPA(Environmental Protection Agency)) and REACH(Registration, Evaluation, Authorization, and Restriction of Chemicals). Regular audits by chemists and toxicologists are essential to refine predictions and maintain ethical and safety standards.

### v5

@@ 7 Open Challenges
While the proposed reference framework offers an adaptable approach for interpreting the HHH principle, it is subject to several limitations that merit consideration. First, the framework does not adequately address fundamental challenges inherent to the HHH principle themselves. For example, it provides limited support for mitigating issues like out-of-distribution (OOD) adversarial attacks (zou2023universal; huang2024obscureprompt; huang2025contextualdistraction), where models encounter inputs significantly deviating from their training data. Such scenarios often require specialized robustness and generalization methods for enhancement, which fall outside the scope of this framework.
@@
Additionally, the framework faces implementation challenges–inherent difficulties associated with multi-objective optimization (MOO). Balancing helpfulness, honesty, and harmlessness requires trade-offs that are complicated by normative assumptions embedded in the optimization objectives. The lack of precise quantification methods for the interaction between these principles complicates achieving a theoretically optimal balance, limiting the framework’s applicability in certain high-stakes or ambiguous scenarios. Following up recent studies propose potential methods for MOO alignment (bai2022training; mukherjee2024multi), more efforts should be made in addressing this challenge.
@@
Moreover, a fundamental question remains: Can human values, ethics, and honesty be embedded into AI systems with sufficient certainty? This challenge is an requirement and pivotal for the proposed framework to function effectively, yet it remains an open problem (sorensen2024roadmap). Human values are inherently dynamic, context-dependent, and often subjective, making it difficult to codify them in a way that guarantees reliable and universally accepted AI behavior (scherrer2024evaluating). Furthermore, ethical principles may conflict in ambiguous scenarios, requiring nuanced decision-making that AI models currently struggle to replicate (e.g., (moore-etal-2024-large; shrivastava2024measuringfreeformdecisionmakinginconsistency)). Without reliable mechanisms to encode these values, the framework risks being constrained by the same uncertainties that challenge broader AI alignment efforts.
@@
Notably, the criticism that HHH principles are merely high-level heuristics and lack operational substance neglects two important facts. First, the original intention of HHH was to anchor alignment efforts in human-centered values, offering interpretable and ethically grounded objectives for training and evaluation. While it is true that these principles are polysemantic and context-sensitive-as are any high-level ethical ideals such as autonomy or respect-this does not diminish their value. Rather, it reinforces the necessity of the adaptive framework we propose, which systematically resolves ambiguities through contextual modeling, priority ordering, and risk assessment. Second, asserting that operational guidance has moved beyond HHH (e.g., to model specs or rule-based reward shaping) overlooks the fact that these newer approaches often implicitly instantiate the HHH values. The 30-40 page model specs (OpenAI2024ModelSpec) and deliberative alignment strategies do not replace HHH; they concretize it. Our contribution lies in formalizing this process: instead of discarding HHH for being abstract, we operationalize it adaptively-preserving its normative clarity while enabling contextual application.
@@ Appendix A Model, Framework and Benchmark Example
HHH (askell2021general)
TrustLLM (huang2024position)
NIST (nist_cybersecurity_framework)
AI Safety and Security (qi2024ai)
Risk Assessment (kapoor2024societal)
Transparency (bommasani2024foundation)
Table 2: Typical principles or frameworks for guiding AI model alignment (excluding government laws or acts, e.g., EU AI Act (eu_ai_act) and AI Bill of Rights (BlueprintAIBill2022)). , , and  indicate different emphasis on trustworthiness or utility.
@@
Llama 2 (touvron2023llama) LLM 2023
LLaVA (liu2024visual) VLM 2023
Qwen (bai2023qwen) LLM 2023
InternLM (team2023internlm) LLM 2023
Llava-med (li2024llava) VLM 2024
Aya model (ustun2024aya) LLM 2024
LLaVA-Plus (liu2025llava) VLM 2024
MiniCPM (hu2024minicpm) LLM 2024
@@
MTBench (zheng2023judging)
MMLU (hendrycks2020measuring)
HumanEval (chen2021evaluating)
Libra-Leaderboard (li2024libra)
DecodingTrust (decodingtrust)
TrustLLM (huang2024trustllm)
HonestLLM (gao2024honestllm)
BeHonest (chern2024behonest)
@@ Appendix B Example of Priority Order
  • Healthcare: Honesty is uttermost critical, as incorrect diagnoses or misleading AI outputs can directly endanger lives. Harmlessness safeguards sensitive medical records, ensuring AI maintains their privacy yet with accurate responses liu2023deid. Helpfulness is desirable but meaningless if accuracy is compromised.
@@ Appendix C Case Study of the 3H principles for developing Chemistry Foundation Models
Alignment Auditing Context-Specific Benchmarks: Design benchmarks that align with the prioritized dimensions. For instance: 1). Measured by the model’s predictive accuracy and relevance to user tasks (guo2023can), 2). Evaluated through safety checks on generated molecular outputs. 3). Assessed by comparing generated outputs with verified scientific data (guo2024can; bushuiev2024massspecgym). Multi-Dimensional Integration: Use complementary benchmarks to capture performance holistically across different dimensions.
Governance Infrastructure Transparency: In the chemistry domain, models must provide interpretable outputs, highlighting key chemical features influencing predictions and annotating data sources such as PubChem (pubchem2021) or ChEMBL (chembl2019). Confidence scores and error margins should accompany predictions to ensure reliability. Governance: It requires strict validation of outputs against experimental data, safeguards against misuse (e.g., toxic compound generation), and compliance with safety standards like EPA(Environmental Protection Agency)) and REACH(Registration, Evaluation, Authorization, and Restriction of Chemicals). Regular audits by chemists and toxicologists are essential to refine predictions and maintain ethical and safety standards.